40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15657     Entry time: Tue Nov 3 09:06:50 2020
Author: gautam 
Type: Update 
Category: General 
Subject: Lab alarm tripped 

I got a call from Calum ~830am today saying some facilities people entered the lab, opened the south entrance door, and tripped the alarm in the process. I came to the lab shortly after and was able to reset the alarm by flipping the switch on the alarm box at the south end entrance to "Alarm OFF". Then, I double checked that the door is closed, and re-enabled the alarm. The particle count at the SP table is not unusually high and the lasers (Oplev HeNe and AUX X) were still on, so doesn't look like any lasting damage was done. The facilities people were apparently wearing laser safety goggles.

ELOG V3.1.3-